Using documented problem solving , calculate the volume of each product and the energy involved in the transformation for the complete combustion of 1,825.0 mL of butane (C4H10) gas and 18.00L of oxygen gas at 760.0 mmHg and 25.0C.
Answer #1

Hi,

Hope you are doing well.


__________________________________________________________________________

Combustion reaction of butane is as follows:

2C4H10(g) + 1302(g) + 8CO2(g) + 10H OLX

No.of moles of butane,

n_{butane}=\frac{1.825\;L}{22.4\;L/mol}=0.08147\;mol

Hence, According to the balanced equation, no.of moles of oxygen required to burn this much butane is given by,

0.08147 mol no2 = - - x 13 = 0.5296 mol

Volume of O2 used, Vo, = no, x 22.4 L/mol = 0.5296 mol x 22.4 L/mol=11.8625 L

No.of moles of CO2 produced,

0.08147 mol - x 8 = 0.32589 mol nco, =

Volume of CO2 produced, Vco, = nco, X 22.4 L/mol = 0.32589 mol x 22.4 L/mol= 7.3 L

No.of moles of H2O produced,

n_{H_{2}O}=\frac{0.08147\;mol}{2}\times 10=0.4074\;mol

Volume of H2O produced, V_{H_{2}O}=n_{H_{2}O}\times22.4\;L/mol=0.4074\;mol\times22.4\;L/mol \mathbf{=9.125\;L}


__________________________________________________________________________

We know that, Enthalpy change for the reaction is given by,

\Delta H_{rxn}=\Delta H_{f}(products)-\Delta H_{f}(reactants)

.

\Rightarrow \Delta H_{rxn}=\left [\Delta H_{f}(CO_{2}_{(g)})+\Delta H_{f}(H_{2}O_{(g)})\right ]- \left [ \Delta H_{f}(C_{4}H_{10}_{(g)})+\Delta H_{f}(O_{2}_{(g)}) \right ]

.

  \Rightarrow \Delta H_{rxn}=\left [\(-393.5\;kJ/mol)\;kJ/mol+(-286.3\;kJ/mol)\right ]- \left [ -126\;kJ/mol+(0\;kJ/mol ) \right ]

\mathbf{\therefore \Delta H_{rxn}=-553.8\;kJ/mol}

This is for 1 mol of butane. For n mole of butane heat released is given by:

\Delta H_{rxn}=-553.8\;kJ/mol\times0.08147\;mol

\mathbf{\therefore \Delta H_{rxn}=-45.12\;kJ}


__________________________________________________________________________

Hope this helped for your studies. Keep learning. Have a good day.

Feel free to clear any doubts at the comment section.


Please don't forget to give a thumps up.

Thank you. :)